OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form. A. y = -2x + 14 B. y = -2x + 5 C. y = -2x + 9 D. y = -2x + 6

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

solve y - 4 = -2(x - 5) for y

jimthompson5910 (jim_thompson5910):

Similar Example y - 7 = -3(x - 8) y - 7 = -3(x) - 3(-8) y - 7 = -3x + 24 y = -3x + 24 + 7 y = -3x + 31

OpenStudy (anonymous):

2x-2(5) 2x+10+(-4) Y=2x+6

OpenStudy (anonymous):

D

jimthompson5910 (jim_thompson5910):

you have y-4 on the left side how do you undo -4 (to move it to the right side)?

jimthompson5910 (jim_thompson5910):

you have to ____ to both sides

OpenStudy (anonymous):

subtract?

OpenStudy (anonymous):

So its A

jimthompson5910 (jim_thompson5910):

what undoes subtraction

OpenStudy (anonymous):

adding

jimthompson5910 (jim_thompson5910):

so you add ____ to both sides

OpenStudy (anonymous):

-4 to both sides thats what i did

jimthompson5910 (jim_thompson5910):

you add -4 to both sides?

jimthompson5910 (jim_thompson5910):

you sure?

OpenStudy (anonymous):

yea i did it how you showed in the example

jimthompson5910 (jim_thompson5910):

in my example, I didn't add -7 to both sides I added 7 to both sides

OpenStudy (anonymous):

the answer is A